# Privately Estimating Black-Box Statistics

### v1

@@ 3.4 Covering Designs
Proof.
The lower bound is due to [Sch64] and the upper bound is due to [ES74, Theorem 13.4]. Both proofs rely on the probabilistic method. We recap both proofs for completeness.
$C(n,m,t)\geq\frac{n}{m}C(n-1,m-1,t-1)\geq\frac{n}{m}\frac{n-1}{m-1}C(n-2,m-2,t-2)\geq\cdots\geq\prod_{i=0}^{t-1}\frac{n-i}{m-i}=\frac{{n\choose t}}{{m\choose t}}$. (3.12)
${\mathbb{P}\left[\not\exists i\in[k_{1}]~~T\subseteq S_{i}\right]}=\prod_{i\in[k_{1}]}\big(1-{\mathbb{P}\left[T\subseteq S_{i}\right]}\big)=\left(1-\frac{{n-t\choose m-t}}{{n\choose m}}\right)^{k_{1}}.$ (3.13)
${\mathbb{E}\left[K_{2}\right]}={n\choose t}{\mathbb{P}\left[\not\exists i\in[k_{1}]~~T\subseteq S_{i}\right]}={n\choose t}\left(1-\frac{{n-t\choose m-t}}{{n\choose m}}\right)^{k_{1}},$ (3.14)
$C(n,m,t)$ $\leq k_{1}+\left\lfloor{n\choose t}\left(1-\frac{{n-t\choose m-t}}{{n\choose m}}\right)^{k_{1}}\right\rfloor$ (3.15)
$C(n,m,t)\leq\left\lceil\frac{{n\choose t}}{{m\choose t}}\log{m\choose t}\right\rceil+\left\lfloor\frac{{n\choose t}}{{m\choose t}}\right\rfloor<\frac{{n\choose t}}{{m\choose t}}\left(\log{m\choose t}+1\right)+1,$ (3.16)
which is the desired result.777[ES74] state the result without the +1+1 at the end, but it is not clear to us how they obtain this result. Their proof ignores the need to round $k_{1}$ to an integer and they leave fixing this issue as an exercise.
$C(n,m,t)\geq\left\lceil\frac{n}{m}C(n-1,m-1,t-1)\right\rceil$ for all $n\geq m\geq t\geq 1$. Equation 3.12 then becomes
$C(n,m,t)\geq\left\lceil\frac{n}{m}C(n-1,m-1,t-1)\right\rceil\geq\left\lceil\frac{n}{m}\left\lceil\frac{n-1}{m-1}\left\lceil\frac{n-2}{m-2}\left\lceil\cdots\left\lceil\frac{n-t+1}{m-t+1}\right\rceil\cdots\right\rceil\right\rceil\right\rceil\right\rceil$ (3.17)
@@
$\left(\frac{n}{m}\right)^{t}\leq\left(\frac{2n-t+1}{2m-t+1}\right)^{t}\leq C(n,m,t)$ $<\left(\frac{n}{m}\cdot\frac{n-t+1}{m-t+1}\right)^{t/2}\cdot\min\left\{\begin{array}[]{c}1+t+t\log(m/t),\\ 1+m\log 2,\\ 1+t\log m\end{array}\right\}+1$ (3.21)
$\leq\left(\frac{n-t+1}{m-t+1}\right)^{t}\cdot\min\{1+m,1+t\log m\}+1.$ (3.22)
Proof.
Proposition 3.7 states that
$\frac{{n\choose t}}{{m\choose t}}\leq C(n,m,t)\leq\frac{{n\choose t}}{{m\choose t}}\left(1+\log{m\choose t}\right)+1.$ (3.23)
We have ${m\choose t}\leq\min\{(em/t)^{t},2^{m},m^{t}\}$. It only remains to bound
$\frac{{n\choose t}}{{m\choose t}}=\prod_{i=0}^{t-1}\frac{n-i}{m-i}=\exp\left(\sum_{i=0}^{t-1}f(i)\right),$ (3.24)
where we define $f(x)=\log\left(\frac{n-x}{m-x}\right)=\log(n-x)-\log(m-x)$. For $0\leq x<m\leq n$, we have $f^{\prime}(x)=\frac{-1}{n-x}-\frac{-1}{m-x}=\frac{n-m}{(m-x)(n-x)}\geq 0$ and $f^{\prime\prime}(x)=\frac{-1}{(n-x)^{2}}-\frac{-1}{(m-x)^{2}}=\frac{(n-x)^{2}-(m-x)^{2}}{(n-x)^{2}(m-x)^{2}}=\frac{(n-m)((n-x)+(m-x))}{(n-x)^{2}(m-x)^{2}}\geq 0$. Thus $f$ is an increasing and convex function on $[0,m)$. Using $\log(\frac{n}{m})=f(0)\leq f(i)\leq f(t-1)=\log(\frac{n-t+1}{m-t+1})$ gives
$\left(\frac{n}{m}\right)^{t}\leq\frac{{n\choose t}}{{m\choose t}}=\exp\left(\sum_{i=0}^{t-1}f(i)\right)\leq\left(\frac{n-t+1}{m-t+1}\right)^{t}.$ (3.25)
Jensen’s inequality gives
$f\left(\frac{t-1}{2}\right)=f\left(\frac{1}{t}\sum_{i=0}^{t-1}i\right)\leq\frac{1}{t}\sum_{i=0}^{t-1}f(i)\leq\frac{1}{t}\sum_{i=0}^{t-1}\frac{(t-1-i)f(0)+if(t-1)}{t-1}=\frac{f(0)+f(t-1)}{2},$ (3.26)
which rearranges to
$\left(\frac{n-\frac{t-1}{2}}{m-\frac{t-1}{2}}\right)^{t}\leq\frac{{n\choose t}}{{m\choose t}}=\exp\left(\sum_{i=0}^{t-1}f(i)\right)\leq\left(\frac{n}{m}\cdot\frac{n-t+1}{m-t+1}\right)^{t/2},$ (3.27)
as required. ∎
@@
Equation 3.30 follows by partitioning $\ell n]$ into $n$ chunks of size $\ell$ and taking a $(n,m,t)$-covering design and applying it to the chunks instead of individual points. Combining Equations 3.29 and 3.30 gives
@@
where $\ell=\frac{n}{t+1}$ is assumed to be an integer.
@@ 4 Our Algorithm
where we set $\max\emptyset=\min\mathcal{Y}$ to cover the corner case. The requirement $|x^{\prime}_{[n]\setminus S_{i}}|=n-|S_{i}|$ is equivalent to requiring that the are no null values in $x^{\prime}_{[n]\setminus S_{i}}$ – i.e., $\forall j\in[n]\setminus S_{i}~~x^{\prime}_{j}\neq\bot$. (Recall how size is defined in Equation 3.1.)
@@ 5 Lower Bound
Now we prove our lower bound which shows the near-optimality of our upper bound.
@@
Let $M^{f}:\mathbb{Z}^{n}\to\mathbb{Z}$ be a randomised algorithm that makes at most $k$ queries to an oracle $f:\mathbb{Z}^{*}\to\mathbb{Z}$. Let $\gamma>0$ be fixed. Suppose that, for every oracle $f$, the algorithm $M^{f}$ satisfies $(\varepsilon,\delta)$-differential privacy and the following. Let $\mathcal{D}$ be an arbitrary distribution on $\mathbb{Z}$ and let $\nu\in\mathbb{Z}$. If ${\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[f(X)=\nu\right]}\geq 1-\gamma$, then ${\underset{X\leftarrow\mathcal{D}^{n}}{\mathbb{P}}\left[|M^{f}(X)-\nu|\leq 1\right]}\geq 1/2$. Then we must have
$k\geq\frac{{n\choose t}}{{m\choose t}}\cdot\left(\frac{1}{2}e^{-2t\varepsilon}-\frac{\delta}{e^{\varepsilon}-1}\right)$ (5.1)
for all integers $t\geq 0$.
Theorem 1.2 in the introduction is attained by substituting either $t=1/\varepsilon$ or $t=\log(1/\delta)/200\varepsilon$ into Theorem 5.1.
We begin by giving some intuition for the lower bound: Consider the function $f:\mathbb{Z}^{*}\to\mathbb{Z}$ being the $\max$ function and consider the data distribution $\mathcal{D}$ to be a point mass on 0. Clearly, $f$ always evaluates to 0 on input from this distribution and thus our differentially private algorithm, given $n$ samples from this distribution, should output something near 0 with high probability. Assume, for now, that $M$ is restricted to evaluating $f$ on subsets of its input of the appropriate size. Now suppose $t\approx\frac{1}{\varepsilon}\log(1/\delta)$ out of the $n$ samples are corrupted to be $1$ instead of 0. If each subset that $M$ evaluates $f$ on includes at least one corrupted sample, then each evaluation will return something near $1$ and so $M$ should output $1$ with high probability. This sets up a contradiction with group privacy, since we have two inputs that only differ by $t$ replacements on which the outputs are very different. The only way to avoid the contradiction is for the algorithm to query enough sets so that at least one of them doesn’t contain any corrupted samples – in a sense, the subsets queried must form a covering design – and that’s the lower bound.
@@
Let $\ell$ be a sufficiently large integer. [^9] We will eventually take $\ell\to\infty$. Let $S\subseteq[\ell^{2}]$ be uniformly random with size $|S|=\ell$. Define $\mathcal{D}_{S}$ to be the uniform distribution on $S$. Let $\nu\in\mathbb{Z}\setminus\{0\}$ be arbitrary. Let $f_{S,\nu,n-m}:\mathbb{Z}^{*}\to\mathbb{Z}$ be defined as
@@
assuming $\ell\geq n^{2}/2\gamma$. Hence, by assumption, $M^{f_{S,\nu,n-m}}(X)\in[\nu-1,\nu+1]$ with probability at least $1/2$ whenever $X\leftarrow\mathcal{D}_{S}^{n}$.
@@
Now we take a Bayesian perspective. Conditioned on $\widetilde{X}$, the set $S$ and the output $Y$ are independent. We can consider a fixed $Y$; and we can assume |Y|=nm|Y|=n-m and $\forall i\neq j~Y_{i}\neq Y_{j}$. Given $\widetilde{X}$ (with no collisions – i.e., $|\widetilde{X}|=n$), all we know about $S$ is that its intersection with $\widetilde{X}$ has size exactly $|S\cap\widetilde{X}|=n-t$; otherwise $S$ is uniformly random with size $|S|=\ell$. Now we can directly calculate the probability that $\forall i~Y_{i}\in S$ based on the intersection between $Y$ and $\widetilde{X}$, namely
@@
It follows from Claim 5.2 and a union bound that if $M^{f_{S,\nu,n-m}}$ makes at most $k$ queries to the oracle $f_{S,\nu,n-m}$, then the probability that the oracle ever returns $\nu$ is at most $k\cdot\left(\frac{{m\choose t}}{{n\choose t}}+\frac{n^{2}}{2\ell}\right)$. Hence we can almost simulate $M^{f_{S,\nu,n-m}}$ by running $M$ with an oracle that always returns 0. Namely, for all $V\subseteq\mathbb{Z}$,
@@
Note that the right hand side of Equation 5.11 does not depend on $\nu$. Now we pick an arbitrary $\nu\in\mathbb{Z}$ such that
${\underset{S,\widetilde{X}\leftarrow\mathcal{C}_{S,n,t}}{\mathbb{P}}\left[M^{0}(\widetilde{X})\in[\nu-1,\nu+1]\right]}\leq\frac{1}{\ell}.$ (5.12)
@@
${\underset{X\leftarrow\mathcal{D}_{S}^{n}}{\mathbb{P}}\left[M^{f_{S,\nu,n-m}}(\widetilde{X})\in V\right]}\leq e^{2t\varepsilon}{\underset{\widetilde{X}\leftarrow\mathcal{C}_{S,n,t}}{\mathbb{P}}\left[M^{f_{S,\nu,n-m}}(\widetilde{X})\in V\right]}+\frac{e^{2t\varepsilon}-1}{e^{\varepsilon}-1}\delta.$ (5.13)
@@
${\underset{X\leftarrow\mathcal{D}_{S}^{n}}{\mathbb{P}}\left[M^{f_{S,\nu,n-m}}(\widetilde{X})\in[\nu-1,\nu+1]\right]}\geq\frac{1}{2}.$ (5.14)
@@
$\frac{1}{2}\leq e^{2t\varepsilon}\left(\frac{1}{\ell}+k\cdot\left(\frac{{m\choose t}}{{n\choose t}}+\frac{n^{2}}{2\ell}\right)\right)+\frac{e^{2t\varepsilon}-1}{e^{\varepsilon}-1}\delta.$ (5.15)
@@
$\frac{1}{2}\leq e^{2t\varepsilon}\cdot k\cdot\frac{{m\choose t}}{{n\choose t}}+\frac{e^{2t\varepsilon}-1}{e^{\varepsilon}-1}\delta\leq e^{2t\varepsilon}\cdot\left(k\cdot\frac{{m\choose t}}{{n\choose t}}+\frac{\delta}{e^{\varepsilon}-1}\right),$ (5.16)
@@
$k\geq\frac{{n\choose t}}{{m\choose t}}\cdot\left(\frac{1}{2}e^{-2t\varepsilon}-\frac{\delta}{e^{\varepsilon}-1}\right),$ (5.17)

### v2

@@ 3.4 Covering Designs
The lower bound is due to [Sch64] and the upper bound is due to [ES74, Theorem 13.4]. Both proofs rely on the probabilistic method.
@@
$\left(\frac{n}{m}\right)^{t}\leq C(n,m,t)<\left(\frac{ne}{m}\right)^{t}\cdot\min\{1+m\log 2,1+t\log m\}+1.$ (3.12)
For completeness, we provide proofs of Proposition 3.7 and Corollary 3.8 in Appendix B.
@@
Equation 3.15 follows by partitioning $\ell n]$ into $n$ chunks of size $\ell$ and taking a $(n,m,t)$-covering design and applying it to the chunks instead of individual points. Combining Equations 3.14 and 3.15 gives
@@
where $\ell=\frac{n}{t+1}$ is assumed to be an integer. Finally, note that we can largely ignore integer divisibility issues due to monotonicity – i.e., $C(n-i,m+j,t)\leq C(n,m,t)$ when $n\geq n-i\geq m+j\geq m\geq t\geq 0$.
@@ 4 Our Algorithm
where we set $\max\emptyset=\min\mathcal{Y}$ to cover the corner case. The requirement $|x^{\prime}_{[n]\setminus S_{i}}|=n-|S_{i}|$ is equivalent to requiring that there are no null values in $x^{\prime}_{[n]\setminus S_{i}}$ – i.e., $\forall j\in[n]\setminus S_{i}~~x^{\prime}_{j}\neq\bot$. (Recall how size is defined in Equation 3.1.)
@@
Note that the proof of Theorem 1.1 uses a union bound over the $k$ evaluations of the function $f$. This bound may be overly pessimistic, since the events are not independent. As a simple thought experiment, suppose each data point is corrupted with probability $\gamma$ and the function evaluation fails only if one of the data points is corrupted. In this case, it suffices to take a union bound over the $n$ data points, rather than over the $k$ evaluations. Theorem 4.3 is stated differently than Theorem 1.1 so that we can use a different analysis, rather than a union bound.
@@ 4.1 Pure & Concentrated DP Variants
4.2 Example Applications
We now consider some example applications of our algorithm. We consider the mean and the maximum. Obviously, there is no need to treat these simple functions as black boxes and tailored algorithms would be more efficient than our method. However, it is instructive to see how well our method works for these examples where we know what outcome to expect.
Gaussian Mean Estimation: Differentially private mean estimation is well-studied [BD14, KV18, BS19, KLSU19, BDKU20, KSU20, WXDX20, DFMBG20, HLY21, CWZ21, BGSUZ21, LKKO21, HKM22, LKO22, KLZ22, KMV22, TCKMS22, AKTVZ23, KDH23, AUZ23, BHS23, HKMN23, ZLSLWL24, ALNP24, Sin24, DJYZZ24, CMC24, YRZ24, KMRSSU25, AKMMSU25, ZZ25, DG26]. We can apply our algorithm to the function $f:(\mathbb{R}\cup\{\bot\})^{*}\to(\mathbb{R}\cup\{\bot\})$ given by $f(x)=\frac{\sum_{i:x_{i}\neq\bot}x_{i}}{\sum_{i:x_{i}\neq\bot}1}$ and compare with known results.
The accuracy achievable for private mean estimation depends on the data distribution [KSU20]. Namely, it depends on how heavy-tailed the data distribution is. For simplicity we consider Gaussian data $\mathcal{D}=\mathcal{N}(\mu,\sigma^{2})$. This has very light tails, which makes it amenable to simple differentially private algorithms.
Corollary 4.6 (Theorem 1.1 applied to Gaussian mean estimation).
Let $\mathcal{Y}\subseteq\mathbb{R}$ be finite. Let $\varepsilon,\delta,\hat{\beta}>0$ and let $t=\frac{1}{\varepsilon}\log(1/\delta)\exp(O(\log^{*}|\mathcal{Y}|))$ as in Theorem 1.1. Let $n\geq m\geq t$. Then there exists a $(\varepsilon,\delta)$-differentially private $M:\mathbb{R}^{*}\to\mathcal{Y}$ with the following property. Let $\mu\in\mathbb{R}$ and $\sigma>0$. If $X\leftarrow\mathcal{N}(\mu,\sigma^{2})^{n}$ consists of $n$ i.i.d. samples from a Gaussian with mean $\mu$ and variance $\sigma^{2}$, then
${\underset{X\leftarrow\mathcal{N}(\mu,\sigma^{2})^{n}}{\mathbb{P}}\left[|M(X)-\mu|\leq O\left(\sigma\sqrt{\frac{t\log(n/m)+t+\log(\log(m)/\hat{\beta})}{n-m}}\right)+\tilde{\mu}_{\mathcal{Y}}\right]}\geq 1-\hat{\beta}$, (4.11)
where $\tilde{\mu}_{\mathcal{Y}}:=\min_{\nu\in\mathcal{Y}}|\mu-\nu|$ is the discretization error from rounding the output to $\mathcal{Y}$.
Setting $m=n/2$ in Equation 4.11 gives
${\underset{X\leftarrow\mathcal{N}(\mu,\sigma^{2})^{n}}{\mathbb{P}}\left[|M(X)-\mu|\leq O\left(\sigma\sqrt{\frac{\frac{1}{\varepsilon}\log(1/\delta)\exp(O(\log^{*}|\mathcal{Y}|))+\log(\log(n)/\hat{\beta})}{n}}\right)+\tilde{\mu}_{\mathcal{Y}}\right]}\geq 1-\hat{\beta}$. (4.12)
The accuracy guarantee of Corollary 4.6 is, unfortunately, suboptimal. Namely, ignoring the other parameters, the optimal accuracy bound [BS19] scales as $\alpha\leq O\left(\frac{\sigma}{\sqrt{n}}+\frac{\sigma}{\varepsilon n}\right)$, whereas Equation 4.12 gives $\alpha\leq O\left(\frac{\sigma}{\sqrt{\varepsilon n}}\right)$.
Proof.
Define $f:(\mathbb{R}\cup\{\bot\})^{*}\to\mathcal{Y}$ to be the closest point in $\mathcal{Y}$ to the mean of the dataset. In symbols, $f(x)=\operatorname*{arg\,min}_{\nu\in\mathcal{Y}}\left|\nu-\frac{\sum_{i:x_{i}\neq\bot}x_{i}}{\sum_{i:x_{i}\neq\bot}1}\right|$ (with $f(x)$ being defined to take an arbitrary fixed value in $\mathcal{Y}$ in the case where $\sum_{i:x_{i}\neq\bot}1=0$). Now we apply Theorem 1.1 to this function to obtain the mechanism $M=M^{f}$. We immediately have the required differential privacy guarantee. It only remains to analyze the accuracy. Theorem 1.1 promises us that
 $\text{ if }{\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[|f(X)-\nu|\leq\alpha\right]}\geq 1-\beta,\text{ then }{\underset{X\leftarrow\mathcal{D}^{n}}{\mathbb{P}}\left[|M^{f}(X)-\nu|\leq\alpha\right]}\geq 1-k\beta,$ (4.13)
where $k=C(n,m,t)\leq O\left(\left(\frac{ne}{m}\right)^{t}\cdot t\log m\right)$. Since $f(X)$ is a mean of nmn-m independent samples from $\mathcal{D}=\mathcal{N}(\mu,\sigma^{2})$, we know its distribution: For all $\beta>0$, we have
${\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[\left|\overline{X}-\mu\right|\leq\sigma\sqrt{\frac{2\log(1/\beta)}{n-m}}\right]}\geq 1-\beta,$ (4.14)
where $\overline{X}:=\frac{1}{n-m}\sum_{i}X_{i}$. Hence
${\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[|f(X)-\mu|\leq 2\sigma\sqrt{\frac{2\log(1/\beta)}{n-m}}+\tilde{\mu}_{\mathcal{Y}}\right]}\geq 1-\beta.$ (4.15)
Equation 4.15 has a multiplicative factor of $2$ and an additive factor of $\tilde{\mu}_{\mathcal{Y}}$ compared to the plain Gaussian tail bound in Equation 4.14. This is because $f$ rounds its output to $\mathcal{Y}$. In particular, if $\mu^{*}_{\mathcal{Y}}=\operatorname*{arg\,min}_{\nu\in\mathcal{Y}}|\nu-\mu|$, we have
$|f(X)-\mu|\leq|f(X)-\overline{X}|+|\overline{X}-\mu|\leq|\mu^{*}_{\mathcal{Y}}-\overline{X}|+|\overline{X}-\mu|\leq|\mu^{*}_{\mathcal{Y}}-\mu|+2|\overline{X}-\mu|=2|\overline{X}-\mu|+\tilde{\mu}_{\mathcal{Y}}.$ (4.16)
The first and last inequalities follow from the triangle inequality and the inequality $|f(X)-\overline{X}|\leq|\mu^{*}_{\mathcal{Y}}-\overline{X}|$ follows from the fact that, by definition, $f(X)$ is the closest point in $\mathcal{Y}$ to $\overline{X}$ and the fact that $\mu^{*}_{\mathcal{Y}}\in\mathcal{Y}$.
Now we substitute Equation 4.15 into Equation 4.13 and set $\beta=\hat{\beta}/k$, so that the final accuracy guarantee has the desired failure probability $\hat{\beta}=k\beta$. This gives the accuracy parameter
$\alpha=2\sigma\sqrt{\frac{2\log(k/\hat{\beta})}{n-m}}+\tilde{\mu}_{\mathcal{Y}}=O\left(\sigma\sqrt{\frac{t\log(n/m)+t+\log(\log(m)/\hat{\beta})}{n-m}}\right)+\tilde{\mu}_{\mathcal{Y}},$ (4.17)
as required. ∎
Figure 1: Accuracy bounds for applying Algorithm 4.1 to $f=\max$ with data $X$ that is i.i.d. uniform in $[0,1]$. The bounds shown are lower bounds and the “ideal” answer is $1$, so higher is better. The vertical axis $1-\alpha$ represents the output value of our algorithm – i.e., ${\underset{X\leftarrow[0,1]^{n}}{\mathbb{P}}\left[M^{f}(X)\geq 1-\alpha\right]}\geq 1-\beta$. The parameter $m$ on the horizontal axis represents the number of elements discarded for privacy and ranges from $t$ to n1n-1. Solid lines represent the union bound given in Theorem 1.1 and dotted lines represent an alternative bound using the order statistics. Blue lines (lowest) represent a 99% confidence bound (i.e., $\beta=0.01$); green lines are 95% confidence ($\beta=0.05$); red lines are 90% confidence ($\beta=0.1$); and yellow lines (highest) are 50% confidence (i.e., median / $\beta=0.5$).
Maximum: Consider the function $f:(\mathbb{R}\cup\{\bot\})^{*}\to(\mathbb{R}\cup\{\bot\})$ given by $f(x)=\max\{x_{i}:x_{i}\neq\bot\}$ (where $\max\emptyset=\bot$). [^8] In other words, $\bot=$-$\infty$. This function has infinite sensitivity and even infinite local sensitivity, so it’s a tough example to apply our algorithm to.
For simplicity, we will ignore the need for the output space $\mathcal{Y}$ to be finite. [^9] In Figure 1, we simply set the parameter $t$ to some fixed value and thus ignore the fact that $t$ depends on the size of the output space $\mathcal{Y}$. Of course, we could discretize the output space, but that would increase the complexity of the analysis without adding any insight.
Now consider data $x$ that is drawn i.i.d. from the uniform distribution on $[0,1]$. In this case, the “ideal” output is $1$, since that is what the output converges to as the sample size increases. Thus the measure of accuracy for our algorithm is how close the output is to $1$.
For a sample of size nmn-m, the maximum follows a $\mathsf{Beta}(n-m,1)$ distribution. This has mean $1-\frac{1}{n-m+1}$. And the cumulative distribution function is given by
${\underset{X_{1},\cdots,X_{n-m}\leftarrow[0,1]}{\mathbb{P}}\left[\max\{X_{1},\cdots,X_{n-m}\}\geq 1-\alpha\right]}=1-(1-\alpha)^{n-m}=1-\beta.$ (4.18)
We can plug this into Theorem 1.1 to get the statistical accuracy guarantee that
${\underset{X_{1},\cdots,X_{n}\leftarrow[0,1]}{\mathbb{P}}\left[M^{f}(X)\geq 1-\alpha\right]}\geq 1-k\beta=1-k(1-\alpha)^{n-m},$ (4.19)
[…]
As discussed on page 4.3, Theorem 1.1 applies a union bound and this bound may be loose. Thus we also consider an alternative bound. Specifically, Theorem 4.3 gives a bound of the form $M(X)\geq\min\{f(X_{[n]\setminus S_{i}}):i\in[k]\}$. Instead of bounding this with a union bound over $i\in[k]$, we can use the fact that each set $n]\setminus S_{i}$ is of size nmn-m and thus $f(X_{[n]\setminus S_{i}})=\max\{X_{j}:j\in[n]\setminus S_{i}\}\geq X_{(n-m)}$, where $X_{(1)}\leq X_{(2)}\leq X_{(3)}\leq\cdots\leq X_{(n)}$ denotes the order statistics of the sample $X$. In particular, $X_{(n-m)}$ follows a $\mathsf{Beta}(n-m,m+1)$ distribution with mean $1-\frac{m+1}{n+1}$. Figure 1 also plots the bound $M(X)\geq X_{(n-m)}$.
In Figure 1 we see several things: First, the solid lines are relatively flat for most of the range, which shows that the bound of Theorem 1.1 is relatively insensitive to the choice of $m$ in this range. This is because two effects cancel out – decreasing $m$ improves the accuracy of each subsample of size nmn-m, but it increases the number of subsamples $k$ that we must union bound over. However, as $m$ approaches $n$, we do see a significant drop in accuracy as the sample size here shrinks. Second, the union bound is indeed not tight, as illustrated by the dotted lines giving a better bound for small values of $m$. Small $m$ is the parameter regime where the subsamples are highly overlapping and thus far from independent. Third, the three different colored lines, which represent different probability bounds, are very close. This indicates that the outcomes (or, at least, our bounds on those outcomes) are highly concentrated.
@@ 5 Lower Bound
Now we prove our lower bound which shows the near-optimality of our upper bound. Figure 2 shows how this bound looks numerically.
Figure 2: Comparing the query complexity lower bound from Theorem 5.1 (taking a max over $t$) with the upper bound from Theorem 4.4 (which satisfies pure DP). The parameters are chosen according to Theorem 4.4 so that on the left $t=2\left\lceil\frac{2}{\varepsilon}\log\left(\frac{|\mathcal{Y}|}{\beta}\right)\right\rceil=100$ and on the right $t=500$, which allows comparison with Figure 1.
@@
Let $\mathcal{Y}\subset\mathbb{Z}$ be finite with at least $2$ elements. Let $M^{f}:\mathbb{Z}^{n}\to\mathcal{Y}$ be a randomised algorithm that makes at most $k$ queries to an oracle $f:\mathbb{Z}^{*}\to\mathcal{Y}$. Let $\gamma>0$ be fixed. Suppose that, for every oracle $f$, the algorithm $M^{f}$ satisfies $(\varepsilon,\delta)$-differential privacy and the following. Let $\mathcal{D}$ be an arbitrary distribution on $\mathbb{Z}$ and let $\nu\in\mathcal{Y}$. If ${\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[f(X)=\nu\right]}\geq 1-\gamma$, then ${\underset{X\leftarrow\mathcal{D}^{n}}{\mathbb{P}}\left[M^{f}(X)=\nu\right]}\geq 1/2$. Then we must have
$k\geq\frac{{n\choose t}}{{m\choose t}}\cdot\left(\frac{1}{2}e^{-2t\varepsilon}-\frac{\delta}{e^{\varepsilon}-1}-\frac{1}{|\mathcal{Y}|-1}\right)$ (5.1)
for all integers $t\in[m]$.
Theorem 1.2 in the introduction is attained by substituting either $t=1/\varepsilon$ or $t=\log(1/\delta)/200\varepsilon$ into Theorem 5.1 and taking $\mathcal{Y}$ to be a sufficiently large set of integers.
We begin by giving some intuition for the lower bound: Consider $\mathcal{Y}=\{0,1\}$ and the function $f:\{0,1\}^{*}\to\{0,1\}$ being the $\max$ function and consider the data distribution $\mathcal{D}$ to be a point mass on 0. Clearly, $f$ always evaluates to 0 on input from this distribution and thus our differentially private algorithm, given $n$ samples from this distribution, should output 0 with high probability. Assume, for now, that $M$ is restricted to evaluating $f$ on subsets of its input of the appropriate size. Now suppose $t\approx\frac{1}{\varepsilon}\log(1/\delta)$ out of the $n$ samples are corrupted to be $1$ instead of 0. If each subset that $M$ evaluates $f$ on includes at least one corrupted sample, then each evaluation will return $1$ and so $M$ should output $1$ with high probability. This sets up a contradiction with group privacy, since we have two inputs that only differ by $t$ replacements on which the outputs are very different. The only way to avoid the contradiction is for the algorithm to query enough sets so that at least one of them doesn’t contain any corrupted samples – in a sense, the subsets queried must form a covering design – and that’s the lower bound.
@@
Let $\ell$ be a sufficiently large integer. [^10] We will eventually take $\ell\to\infty$. Let $S\subseteq[\ell^{2}]$ be uniformly random with size $|S|=\ell$. Define $\mathcal{D}_{S}$ to be the uniform distribution on $S$. Without loss of generality, assume $0\in\mathcal{Y}$. Let $\nu\in\mathcal{Y}\setminus\{0\}$ be arbitrary. Let $f_{S,\nu,n-m}:\mathbb{Z}^{*}\to\mathcal{Y}$ be defined as
@@
assuming $\ell\geq n^{2}/2\gamma$. Hence, by assumption, $M^{f_{S,\nu,n-m}}(X)=\nu$ with probability at least $1/2$ whenever $X\leftarrow\mathcal{D}_{S}^{n}$.
@@
Now we take a Bayesian perspective. Conditioned on $\widetilde{X}$, the set $S$ and the output $Y$ are independent. We want to upper bound the probability of the event that $Y$ consists of nmn-m distinct points all in $S$. We can consider a fixed $Y$; and we can assume |Y|=nm|Y|=n-m and $\forall i\neq j~Y_{i}\neq Y_{j}$ (since otherwise the probability of the event is zero). Given $\widetilde{X}$ (with no collisions – i.e., $|\widetilde{X}|=n$), all we know about $S$ is that its intersection with $\widetilde{X}$ has size exactly $|S\cap\widetilde{X}|=n-t$; otherwise $S$ is uniformly random with size $|S|=\ell$. Now we can directly calculate the probability that $\forall i~Y_{i}\in S$ based on the intersection between $Y$ and $\widetilde{X}$, namely
@@
It follows from Claim 5.2 and a union bound that, if $M^{f_{S,\nu,n-m}}$ makes at most $k$ queries to the oracle $f_{S,\nu,n-m}$, then the probability that the oracle ever returns something other than 0 is at most $k\cdot\left(\frac{{m\choose t}}{{n\choose t}}+\frac{n^{2}}{2\ell}\right)$. Hence we can almost simulate $M^{f_{S,\nu,n-m}}$ by running $M$ with an oracle that always returns 0. Namely, for all $V\subseteq\mathcal{Y}$,
@@
Note that the right hand side of Equation 5.11 does not depend on $\nu$. Now we pick an arbitrary $\nu\in\mathcal{Y}\setminus\{0\}$ such that
${\underset{S,\widetilde{X}\leftarrow\mathcal{C}_{S,n,t}}{\mathbb{P}}\left[M^{0}(\widetilde{X})=\nu\right]}\leq\frac{1}{|\mathcal{Y}\setminus\{0\}|}.$ (5.12)
@@
${\underset{X\leftarrow\mathcal{D}_{S}^{n}}{\mathbb{P}}\left[M^{f_{S,\nu,n-m}}({X})\in V\right]}\leq e^{2t\varepsilon}{\underset{\widetilde{X}\leftarrow\mathcal{C}_{S,n,t}}{\mathbb{P}}\left[M^{f_{S,\nu,n-m}}(\widetilde{X})\in V\right]}+\frac{e^{2t\varepsilon}-1}{e^{\varepsilon}-1}\delta.$ (5.13)
@@
${\underset{X\leftarrow\mathcal{D}_{S}^{n}}{\mathbb{P}}\left[M^{f_{S,\nu,n-m}}({X})=\nu\right]}\geq\frac{1}{2}.$ (5.14)
@@
$\frac{1}{2}\leq e^{2t\varepsilon}\left(\frac{1}{|\mathcal{Y}\setminus\{0\}|}+k\cdot\left(\frac{{m\choose t}}{{n\choose t}}+\frac{n^{2}}{2\ell}\right)\right)+\frac{e^{2t\varepsilon}-1}{e^{\varepsilon}-1}\delta.$ (5.15)
@@
$\frac{1}{2}\leq\frac{e^{2t\varepsilon}}{|\mathcal{Y}\setminus\{0\}|}+e^{2t\varepsilon}\cdot k\cdot\frac{{m\choose t}}{{n\choose t}}+\frac{e^{2t\varepsilon}-1}{e^{\varepsilon}-1}\delta\leq e^{2t\varepsilon}\cdot\left(\frac{1}{|\mathcal{Y}|-1}+k\cdot\frac{{m\choose t}}{{n\choose t}}+\frac{\delta}{e^{\varepsilon}-1}\right),$ (5.16)
@@
$k\geq\frac{{n\choose t}}{{m\choose t}}\cdot\left(\frac{1}{2}e^{-2t\varepsilon}-\frac{\delta}{e^{\varepsilon}-1}-\frac{1}{|\mathcal{Y}|-1}\right),$ (5.17)